# Mirror Descent Policy Optimisation for Robust Constrained Markov Decision Processes

### v2

@@ 5.4 Analysis of Robust Sample-based PMD-PD
$T$ $=\sum_{k=1}^{K}\left(M_{V,k}N_{V,k}+\sum_{t=0}^{t_{k}-1}M_{Q,k}N_{Q,k}+\sum_{t=0}^{t_{k}^{\prime}-1}M_{G,t}N_{G,t}\right)=\tilde{\mathcal{O}}(\epsilon^{-3})\,.$
@@ 5.5 KL-based Mirror Descent in continuous state-action spaces
To generalise the algorithm to continuous state-action spaces, we follow a similar algorithm but replace the tabular approximation with function approximation (see Algorithm 3). A further required change is related to the weighted Bregman divergence term formulated in Eq. 13. While the weighted Bregman divergence is not a proper Bregman divergence over policies, it is over the occupancy distribution, i.e. $B_{d_{\rho}^{\pi,p}}(\pi,\pi^{\prime})=B(d_{\rho}^{\pi,p},d_{\rho}^{\pi^{\prime},p})$, as shown in the pseudo-KL divergence in Lemma 10 of Liu et al., 2021b . To maintain similar theoretical results in continuous state-action spaces, we extend the discrete pseudo-KL divergence of occupancy to a continuous pseudo-KL divergence of occupancy.
@@
5.6 MDPO-Robust-Lag: a practical implementation
To implement the algorithm in practice, we use Mirror Descent Policy Optimisation (MDPO) (Tomar et al.,, 2022). On-policy MDPO optimises the objective
@@ 5.6 MDPO-Robust-Lag: a practical implementation
where in our case the advantage $\hat{A}_{\pi_{\theta_{k}},p_{k}}(s,a)$ is implemented based on the Lagrangian. The use of the importance weight $\frac{\pi_{\theta}}{\pi_{\theta_{k}}}$ is used in MDPO to correct for the deviation from mirror descent theory, where it is assumed that the data at each epoch is generated from $\pi_{\theta}$ rather than $\pi_{\theta_{k}}$. In line with our theory, our implementation performs the critic updates together with policy updates, which deviates from the original MDPO implementation. There still remains some gap of the practical implementation to our theoretical algorithm in that the state occupancy of MDPO is given by the old policy whereas in theory it is based on the current. The implementation of advantage values is based on generalised advantage estimation (Schulman et al.,, 2018) and the critic estimates the value and the constraint-costs based on an MLP architecture with $1+m$ heads. The update of the multipliers is based on the observed constraint-costs in separate samples.
@@
To form a robust-constrained variant of MDPO, we use Lagrangian relaxation similar to the above (i.e. Algorithm 2), which differs from Adversarial RCPG (Bossens,, 2024) in that multiple constraints are considered and that we preserve the structure of the PMD-PD algorithm (Liu et al., 2021b, ) and the use of LTMA similar follows the TMA implementation from Wang et al., (2024), which applies a projected gradient descent based on clipping to the uncertainty set bounds. Due to its similarity to PPO-Lagrangian (or PPO-Lag for short) (Ray et al.,, 2019) and the use of updates in the robust Lagrangian, we name the algorithm MDPO-Robust-Lag. While the MDPO-Robust-Lag is based on traditionally clipping the multiplier, we also implement a version with the augmented Lagrangian as suggested by the theory and shown in Algorithm 2 and  3, called MDPO-Robust-Augmented-Lag. The practical implementation of the KL estimate follows the standard implementation in StableBaselines3, i.e. $D_{\text{KL}}(\pi(\cdot|s),\pi^{\prime}(\cdot|s))\approx\mathbb{E}\left[e^{\log(\pi(a|s))-\log(\pi^{\prime}(a|s))}-1-(\log(\pi(a|s))-\log(\pi^{\prime}(a|s)))\right]$.
Experiments
@@
To demonstrate the use of MDPO for RCMDPs, we assess MDPO-Robust-Lag on three RCMDP domains. We compare MDPO-Robust-Lag to robust-constrained algorithms from related algorithm families, namely Monte Carlo based optimisation (RMCPMD (Wang et al.,, 2024) in particular) and proximal policy optimisation with function approximiation (PPO-Lag (Achiam et al.,, 2017) in particular). All the involved algorithm families (MDPO, MC, and PPO) are included with four variants, namely with and without Lagrangian (indicated by the -Lag suffix) and with and without robustness training (indicated by the R- prefix or the -Robust- infix). Additionally, the MDPO-Lag implementations also have Augmented-Lag variants which use the augmented Lagrangian as proposed in the theory section. Further implementation details of the algorithms included to the study can be found in Appendix E.
@@
After training with the above-mentioned algorithms, the agents are subjected to a test which presents transition dynamics based on the distortion levels setup in Wang et al., (2024). The test procedure follows the same general procedure for all domains. For distortion level $x\in[0,1]$, the agent is subjected to a perturbation such that if the nominal model is $\xi_{c}$ and the uncertainty set varies the parameter in $\underline{\xi},\overline{\xi}]$, the transition dynamics kernel parameter of the test with distortion level $x$ is given for each dimension by either $\xi_{i}=\xi_{c}+x^{2}(\overline{\xi}-\xi_{c})$ or $\xi_{i}=\xi_{c}+x^{2}(\underline{\xi}-\xi_{c})$. We slightly improve the test evaluation by considering all the $\pm$ directions rather than just a single one, amounting to a total number of test evaluations $N=n_{\text{test}}\times 2^{n}$, where $n$ is the dimensionality of the uncertainty set parametrisation and $n_{\text{test}}$ is the number of episodes per perturbation.
@@
To summarise the results, we use the penalised return, a popular summary statistic for robust constrained RL (Mankowitz et al.,, 2020; Bossens,, 2024). The statistic is defined for a maximisation problem as $R_{\text{pen}}=V(\rho)-\lambda_{\text{max}}\sum_{j=1}^{m}\max(0,C_{j}(\rho))$ where $V(\rho)$ denotes the value (negative of the cost) from the starting distribution and $C_{j}(\rho)$ denotes the $j$’th constraint-cost from the starting distribution. In addition, we also formulate the signed penalised return $R_{\text{pen}}^{\pm}=V(\rho)-\sum_{j=1}^{m}\lambda_{\text{max}}C_{j}(\rho)$. We note that the penalised return matches the objective of the constrained methods and the return matches the objective of the unconstrained methods. The signed penalised return might be especially useful for robust problems; as even our extensive testing procedure may not find the worst case in the uncertainty set, and it is also of interest to generalise even beyond the uncertainty set, a solution that has negative cost can be evaluated more positively since it is more likely not to have a positive cost in the worst-case transition kernel. The summary statistics reported included both the mean and the minimum since the minimum indicates the effectiveness in dealing with the minimax problem (the worst-case robustness). However, as a note of caution, the minimum is challenging to establish accurately because even though we test many environments it may still not contain the worst-case environment.
@@ 6.1 Cartpole
To assess our algorithm on an RCMDP, we introduce the robust constrained variant of the well-known Cartpole problem (Barto et al.,, 1983; Brockman et al.,, 2016) by modifying the RMDP from Wang et al., (2024). The problem involves a mechanical simulation of a frictionless cart moving back and forth to maintain a pole, which is attached to the cart as an un-actuated joint, in the upright position. The agent observes $x,\dot{x},\theta,\dot{\theta}$ and takes actions in $\{\text{left},\text{right}\}$, which apply a small force moving either left or right. The agent receives a reward of +1+1 until either the cart goes out of the [2.4,2.4][-2.4,2.4]m bounds or the pole has an angle outside the range $12^{\circ},12^{\circ}]$. In contrast to the traditional problem, the mechanics are not deterministic, and following Wang et al., (2024), transitions dynamics models are multi-variate Gaussians of the form
@@
To incorporate robustness into our setting, we introduce delta-perturbations as in Wang et al., (2024), using an uncertainty set with parametrisation
@@
where $\delta\in[-\kappa_{i},\kappa_{i}]_{i=1}^{n}$. Our parameter settings differ in that we use a larger uncertainty set, with $\kappa_{i}$ being increased fivefold (making the problem more challenging) while the number of time steps is reduced to make the constraint satisfiable. The robust algorithms are trained on this uncertainty set, where the transition dynamics are adjusted based on the transition mirror ascent algorithm of Wang et al., (2024), i.e. by using Monte Carlo and mirror ascent. In the experiment, this amounts to a projected update, where the projection restricts the update to lie within the uncertainty set.
@@ 6.2 Inventory Management
A second domain is the Inventory Management domain from Wang et al., (2024), which involves maintaining an inventory of resources. The resources induce a period-wise cost and the goal is to purchase and sell resources such that minimal cost is incurred over time. To form a constrained variant of this benchmark, we introduce the constraint that the discounted sum of actions should not average to higher than zero (i.e. the rate of selling should not exceed that of purchasing). We use the same radial features and clipped uncertainty set as in the implementation on github https://github.com/JerrisonWang/JMLR-DRPMD. The implementation uses Gaussian parametric transition dynamics
@@
where $\mu_{\zeta_{1}}=(-4,5)$, $\mu_{\zeta_{2}}=(-2,8)$. The uncertainty set is given by $\{\eta:\left\lVert\eta-\eta_{c}\right\rVert_{\infty}\leq\kappa\}$ where $\eta_{c}=(-2,3.5)$. With the exception of the constraint, and the number of steps per episode for the adversary, the setting matches Wang et al., (2024). The constraint at time $t$ is given by
@@
Conclusion
This paper presents mirror descent policy optimisation for robust constrained MDPs (RCMDPs), making use of policy gradient techniques to optimise both the policy and the transition kernel (as an adversary) on the Lagrangian representing a constrained MDP. In the sample-based RCMDP setting, we require $\tilde{\mathcal{O}}(\epsilon^{-3})$ samples for an average regret of at most $\epsilon$, confirming an $\tilde{\mathcal{O}}\left(\frac{1}{T^{1/3}}\right)$ convergence rate. An approximate algorithm for transition mirror ascent is also proposed which is of independent interest for designing adversarial environments in general MDPs. Experiments confirm the performance benefits of mirror descent policy optimisation in practice, obtaining significant improvements in test performance.
@@ Appendix A Bregman divergence and associated policy definitions
Parametrisation Bregman divergence Update rule Direct: $\pi=\theta$ $D_{\text{SE}}(\theta,\theta_{t})$ $\theta_{t+1}\leftarrow\text{proj}_{\Delta(\mathcal{A})}(\theta_{t}-\eta_{t}\nabla_{\theta}V_{\pi,p}(d_{0}))$ Softmax: $\pi(a|s)=\frac{\exp(\theta_{s,a})}{\sum_{a^{\prime}}\exp(\theta_{s,a^{\prime}})}$ $D_{\text{SE}}(\theta,\theta_{t})$ $\theta_{t+1}\leftarrow\theta_{t}-\eta_{t}\nabla_{\theta}V_{\pi,p}(d_{0})$ Softmax: $\pi(a|s)=\frac{\exp(\theta_{s,a})}{\sum_{a^{\prime}}\exp(\theta_{s,a^{\prime}})}$ Occupancy-weighted KL-divergence (Eq. 13) $\pi_{k}^{t+1}(a|s)=\frac{1}{Z_{k}^{t}(s)}(\pi_{k}^{t}(a|s))^{1-\frac{\eta\alpha}{1-\gamma}}e^{\frac{-\eta\mathbf{Q}_{\pi_{k}^{t},p}^{\alpha}(s,a)}{1-\gamma}}$
@@ C.1 Recursion property
Lemma 11 (Recursion property, Lemma 13 in Xiao, (2022)).
Let $\beta\in(0,1)$, and $b>0$. If $\{a_{t}\}_{t\geq 0}$ is a non-negative sequence such that $x_{t+1}\leq\beta a_{t}+b$, then $x_{t}\leq\beta^{t}x_{0}+\frac{b}{1-\gamma}$.

### v4

@@ 5.5 Extension to non-rectangular uncertainty sets
$\frac{1}{K}\sum_{k=1}^{K}\left(\Phi(\pi_{k})-\Phi(\pi^{*})\right)\leq\frac{1}{K}\sum_{k=1}^{K}\left(\mathbf{V}_{\pi_{k},p_{k}}(\rho;\lambda_{k-1})+\mathcal{O}(\epsilon+\delta_{\mathcal{P}})+\Delta_{\lambda_{k}}-\mathbf{V}_{\pi^{*},p_{k};\lambda_{k-1}}(\rho)\right)=\mathcal{O}(\epsilon+\delta_{\mathcal{P}})\,.$
Since $n=\tilde{\mathcal{O}}(1/\epsilon^{2})$ in Approximate TMA, substituting $\sum_{k=1}^{K}t_{k}^{\prime}n=\tilde{\mathcal{O}}(1/\epsilon^{5})$ into Eq. 42, the term dominates the sum, resulting in a $T=\tilde{\mathcal{O}}(1/\epsilon^{5})$ sample complexity. ∎
@@ 5.6 KL-based Mirror Descent in continuous state-action spaces
To generalise the algorithm to continuous state-action spaces, we follow a similar algorithm but replace the tabular approximation with function approximation (see Algorithm 4). A further required change is related to the weighted Bregman divergence term formulated in Eq. 16. While the weighted Bregman divergence is not a proper Bregman divergence over policies, it is over the occupancy distribution, i.e. $B_{d_{\rho}^{\pi,p}}(\pi,\pi^{\prime})=B(d_{\rho}^{\pi,p},d_{\rho}^{\pi^{\prime},p})$, as shown in the pseudo-KL divergence in Lemma 10 of Liu et al. (2021b). To maintain similar theoretical results in continuous state-action spaces, we extend the discrete pseudo-KL divergence of occupancy to a continuous pseudo-KL divergence of occupancy.
@@
5.7 MDPO-Robust-Lagrangian: a practical implementation
To implement the algorithm in practice, we use MDPO (Tomar et al., 2022) as the policy optimiser. On-policy MDPO optimises the objective
@@ 5.7 MDPO-Robust-Lagrangian: a practical implementation
where in our case the advantage $\hat{A}_{\pi_{\theta_{k}},p_{k}}(s,a)$ is implemented based on the Lagrangian. The use of the importance weight $\frac{\pi_{\theta}}{\pi_{\theta_{k}}}$ is used in MDPO to correct for the deviation from mirror descent theory, where it is assumed that the data at each epoch is generated from $\pi_{\theta}$ rather than $\pi_{\theta_{k}}$. In line with our theory, our implementation performs the critic updates together with policy updates, which deviates from the original MDPO implementation. There still remains some gap of the practical implementation to our theoretical algorithm in that the state occupancy of MDPO is given by the old policy whereas in theory it is based on the current. The implementation of advantage values is based on generalised advantage estimation (Schulman et al., 2018) and the critic estimates the value and the constraint-costs based on an MLP architecture with $1+m$ outputs. The update of the multipliers is based on the observed constraint-costs in separate samples.
@@
To form a robust-constrained variant of MDPO, we use Lagrangian relaxation similar to the above (i.e. Algorithm 3). The algorithm considers multiple constraints and preserves the structure of the PMD-PD algorithm (Liu et al., 2021b). The LTMA algorithm implementation follows the TMA implementation from Wang et al. (2024), which applies a projected gradient descent based on clipping to the uncertainty set bounds. Due to its similarity to PPO-Lagrangian (or PPO-Lag for short) (Ray et al., 2019) and the use of updates in the robust Lagrangian, we name the algorithm MDPO-Robust-Lagrangian (or MDPO-Robust-Lag for short). While MDPO-Robust-Lag is based on traditionally clipping the multiplier, we also implement MDPO-Robust-Augmented-Lag, which uses the augmented Lagrangian according to the theory (as summarised in Algorithm 3 and  4). The practical implementation of the KL estimate follows the standard implementation in StableBaselines3, i.e. $D_{\text{KL}}(\pi(\cdot|s),\pi^{\prime}(\cdot|s))\approx\mathbb{E}\left[e^{\log(\pi(a|s))-\log(\pi^{\prime}(a|s))}-1-(\log(\pi(a|s))-\log(\pi^{\prime}(a|s)))\right]$.
6 Experiments
@@
To demonstrate the use of MDPO for RCMDPs, we assess MDPO-Robust-Lag on three RCMDP domains. We compare MDPO-Robust-Lag to robust-constrained algorithms from related algorithm families, namely Monte Carlo based mirror descent (RMCPMD (Wang et al., 2024) in particular) and proximal policy optimisation with function approximation (PPO-Lag (Achiam et al., 2017) in particular). All the involved algorithm families (MDPO, MCPMD, and PPO) are included with four variants, namely with and without Lagrangian (indicated by the -Lag suffix) and with and without robustness training (indicated by the R- prefix or the -Robust- infix). Additionally, the MDPO-Lag implementations also have Augmented-Lag variants which use the augmented Lagrangian as proposed in the theory section. The implementation details of the algorithms included in the study can be found in Appendix E. The code for all experiments can be found at https://github.com/bossdm/MDPO-RCMDP.
@@
After training with the above-mentioned algorithms, the agents are subjected to a test which presents transition dynamics based on the distortion levels setup in Wang et al. (2024). The test procedure follows the same general procedure for all domains. For distortion level $x\in[0,1]$, the agent is subjected to a perturbation such that if the nominal model is $\xi_{c}$ and the uncertainty set varies the parameter in $\underline{\xi},\overline{\xi}]$, the transition dynamics kernel parameter of the test with distortion level $x$ is given for each dimension by either $\xi_{i}=\xi_{c}+x^{2}(\overline{\xi}-\xi_{c})$ or $\xi_{i}=\xi_{c}+x^{2}(\underline{\xi}-\xi_{c})$. We slightly improve the test evaluation by considering all the $\pm$ directions rather than just a single one, amounting to a total number of test evaluations $N=n_{\text{test}}\times 2^{n}$, where $n$ is the dimensionality of the uncertainty set parametrisation and $n_{\text{test}}$ is the number of episodes per perturbation.
@@
To summarise the results, we use the penalised return, a popular summary statistic for robust constrained RL (Mankowitz et al., 2020; Bossens, 2024). The statistic is defined for a maximisation problem as $R_{\text{pen}}=V(\rho)-\lambda_{\text{max}}\sum_{j=1}^{m}\max(0,C_{j}(\rho))$ where $V(\rho)$ denotes the value (negative of the cost) from the starting distribution and $C_{j}(\rho)$ denotes the $j$’th constraint-cost from the starting distribution. In addition, we also formulate the signed penalised return $R_{\text{pen}}^{\pm}=V(\rho)-\sum_{j=1}^{m}\lambda_{\text{max}}C_{j}(\rho)$. We note that the penalised return matches the objective of the constrained methods and the return matches the objective of the unconstrained methods. The signed penalised return might be especially useful for robust problems; as even our extensive testing procedure may not find the worst case in the uncertainty set, and it is also of interest to generalise even beyond the uncertainty set, a solution that has negative cost can be evaluated more positively since it is more likely not to have a positive cost in the worst-case transition kernel. The summary statistics reported included both the mean and the minimum since the minimum indicates the effectiveness in dealing with the minimax problem (the worst-case robustness). However, as a note of caution, the minimum is challenging to establish accurately because even though we test many environments it may still not contain the worst-case environment.
@@ 6.1 Cartpole
To assess our algorithm on an RCMDP, we introduce the robust constrained variant of the well-known Cartpole problem (Barto et al., 1983; Brockman et al., 2016) by modifying the RMDP from Wang et al. (2024). The problem involves a mechanical simulation of a frictionless cart moving back and forth to maintain a pole, which is attached to the cart as an un-actuated joint, in the upright position. The agent observes $x,\dot{x},\theta,\dot{\theta}$ and takes actions in $\{\text{left},\text{right}\}$, which apply a small force moving either left or right. The agent receives a reward of +1+1 until either the cart goes out of the [2.4,2.4][-2.4,2.4]m bounds or the pole has an angle outside the range $12^{\circ},12^{\circ}]$. In contrast to the traditional problem, the mechanics are not deterministic, and following Wang et al. (2024), transitions dynamics models are multi-variate Gaussians of the form
@@
To incorporate robustness into our setting, we introduce delta-perturbations as in Wang et al. (2024), using an uncertainty set with parametrisation
@@
where $\delta\in[-\kappa_{i},\kappa_{i}]_{i=1}^{n}$. Our parameter settings differ in that we use a larger uncertainty set, with $\kappa_{i}$ being increased fivefold (making the problem more challenging) while the number of time steps is reduced to make the constraint satisfiable. The robust algorithms are trained on this uncertainty set, where the transition dynamics are adjusted based on the transition mirror ascent algorithm of Wang et al. (2024), i.e. by using Monte Carlo and mirror ascent. In the experiment, this amounts to a projected update, where the projection restricts the update to lie within the uncertainty set.
@@ 6.2 Inventory Management
A second domain is the Inventory Management domain from Wang et al. (2024), which involves maintaining an inventory of resources. The resources induce a period-wise cost and the goal is to purchase and sell resources such that minimal cost is incurred over time. To form a constrained variant of this benchmark, we introduce the constraint that the discounted sum of actions should not average to higher than zero (i.e. the rate of selling should not exceed that of purchasing). We use the same radial features and clipped uncertainty set as in the implementation on github https://github.com/JerrisonWang/JMLR-DRPMD. The implementation uses Gaussian parametric transition dynamics
@@
where $\mu_{\zeta_{1}}=(-4,5)$, $\mu_{\zeta_{2}}=(-2,8)$. The uncertainty set is given by $\{\eta:\left\lVert\eta-\eta_{c}\right\rVert_{\infty}\leq\kappa\}$ where $\eta_{c}=(-2,3.5)$. With the exception of the constraint, and the number of steps per episode for the adversary, the setting matches Wang et al. (2024). The constraint at time $t$ is given by
@@
7 Conclusion
This paper presents mirror descent policy optimisation for RCMDPs, making use of policy gradient techniques to optimise both the policy and the transition kernel (as an adversary) on the Lagrangian representing a CMDP. In the sample-based rectangular RCMDP setting, we require $\tilde{\mathcal{O}}(\epsilon^{-3})$ samples for an average regret of at most $\epsilon$, confirming an $\tilde{\mathcal{O}}\left(1/T^{1/3}\right)$ convergence rate. Our analysis also covers the case of non-rectangular RCMDPs, where our algorithm provides an $\tilde{\mathcal{O}}\left(1/T^{1/5}\right)$ sample-based rate. An approximate algorithm for transition mirror ascent is also proposed which is of independent interest for designing adversarial environments in general MDPs. Experiments confirm the performance benefits of mirror descent policy optimisation in practice, obtaining significant improvements in test performance.
Acknowledgments
This research/project is supported by the National Research Foundation, Singapore under its National Large Language Models Funding Initiative (AISG Award No: AISG-NMLP-2024-003). Any opinions, findings and conclusions or recommendations expressed in this material are those of the author(s) and do not reflect the views of National Research Foundation, Singapore. This research is supported by the National Research Foundation, Singapore and the Ministry of Digital Development and Information under the AI Visiting Professorship Programme (award number AIVP-2024-004). Any opinions, findings and conclusions or recommendations expressed in this material are those of the author(s) and do not reflect the views of National Research Foundation, Singapore and the Ministry of Digital Development and Information.
@@ Appendix A Bregman divergence and associated policy definitions
Parametrisation Bregman divergence Update rule Direct: $\pi=\theta$ $D_{\text{SE}}(\theta,\theta_{t})$ $\theta_{t+1}\leftarrow\text{proj}_{\Pi}(\theta_{t}-\eta_{t}\nabla_{\theta}V_{\pi,p}(d_{0}))$ Softmax: $\pi(a|s)=\frac{\exp(\theta_{s,a})}{\sum_{a^{\prime}}\exp(\theta_{s,a^{\prime}})}$ $D_{\text{SE}}(\theta,\theta_{t})$ $\theta_{t+1}\leftarrow\theta_{t}-\eta_{t}\nabla_{\theta}V_{\pi,p}(d_{0})$ Softmax: $\pi(a|s)=\frac{\exp(\theta_{s,a})}{\sum_{a^{\prime}}\exp(\theta_{s,a^{\prime}})}$ Occupancy-weighted KL-divergence (Eq. 16) $\pi_{k}^{t+1}(a|s)=\frac{1}{Z_{k}^{t}(s)}(\pi_{k}^{t}(a|s))^{1-\frac{\eta\alpha}{1-\gamma}}e^{\frac{-\eta\mathbf{Q}_{\pi_{k}^{t},p}^{\alpha}(s,a)}{1-\gamma}}$
@@ C.1 Recursion property
Lemma 12 (Recursion property, Lemma 13 in Xiao (2022)).
Let $\beta\in(0,1)$, and $b>0$. If $\{a_{t}\}_{t\geq 0}$ is a non-negative sequence such that $x_{t+1}\leq\beta a_{t}+b$, then $x_{t}\leq\beta^{t}x_{0}+\frac{b}{1-\beta}$.